You are provided the following information: Current ratio 1.6 Net profit margin .1 Acid test ratio...

  1. You are provided the following information:

Current ratio

1.6

Net profit margin

.1

Acid test ratio

1.0

Return on common equity

.20

Debt to equity ratio

2.5

Cost of goods sold

12,000

Long term debt to equity ratio

1.0

Dividends

0

Gross profit margin

.25

Preferred stock

0

What is the value of inventory?
Answer #1

For inventory, we need to calculate various other items as beow:

First we will calculate Sales as per below:

Gross profit margin = Sales - Cost of goods sold / Sales

0.25 = Sales - 12000 / Sales

0.25 * Sales = Sales - 12000

0.25 * Sales - Sales = -12000

-0.75 * Sales = -12000

Sales = 12000 / 0.75 = 16000

Next we will calculate net income as per below:

Net profit margin = Net income / Sales

0.1 = Net income / 16000

Net income = 16000 * 0.1

Net income = 1600

Next we will calculate common equity as per below:

Return on common equity = Net income / Common equity

0.20 = 1600 / Common equity

Common equity = 1600 / 0.2 = 8000

Next we will calculate total debt as per below:

Debt equity ratio = Total debt / Equity

2.5 = Total debt / 8000

Total debt = 8000 * 2.5 = 20000

Next we will calculate long term debt as per below:

Long term debt to equity ratio = Long term debt / Equity

1 = Long term debt / 8000

Long term debt = 8000 * 1 = 8000

Next we will calculate current liabilities as per below:

Total debt = Current liabilities + Long term debt

20000 = Current liabilities + 8000

Current liabilities = 20000 - 8000 = 12000

Next we will calculate current assets as per below:

Current ratio = Current assets / Current liabilities

1.6 = Current assets / 12000

Current assets = 12000 * 1.6 = 19200

Now, we will calculate inventory as per below:

Acid test ratio = Current assets - Inventory / Current liabilities

1 = (19200 - Inventory) / 12000

12000 = 19200 - Inventory

Inventory = 19200 - 12000 = 7200

So, inventory = 7200